OpenStudy (anonymous):

Carrie is moving from Mineral Wells to Lubbock and rented a truck from U-Move-It truck rentals. The cost of a one-day truck rental is given by C(m) = 0.6m + 40 where m is the number of miles driven. If Carrie drives 320 miles, what is the cost of the truck rental?

OpenStudy (anonymous):

@satellite73 could you help me?

OpenStudy (anonymous):

replace \(m\) by 320, compute \(C(m)=0.6\times 320+40\)

OpenStudy (anonymous):

\[C(320)=0.6\times 320+40=192+40=232\]

OpenStudy (anonymous):

and this is the same? here? Katie paid $18,802.80 for a Mustang. If the price paid includes an 8% sales tax, which of the following equations can be used to determine the price of the Mustang before tax? (Let x represent the cost of the Mustang and y represent the total cost after tax.)

OpenStudy (anonymous):

\[18,802.80=x+.08x\] \[18,802.80=(1.08)x\] \[x=18.802.80\div (1.08)\]

OpenStudy (anonymous):

no it is not the same, it is a completely different question

OpenStudy (anonymous):

oh I get it, okay :) thanks

OpenStudy (anonymous):

yw